OpenStudy (anonymous):

Solve this for me please? Test! x - 3 = -14 x =

OpenStudy (anonymous):

Add 3 to both sides.

TheSmartOne (thesmartone):

You have to add 3 to both sides x-3+3=-14+3 x=-11